# Universal activated aging and weak ergodicity breaking in spin and structural glasses

### v4

@@ IX Discussion
Acknowledgements.
We warmly thank Wenyue Fan, Gang Huang, Wencheng Ji, Walter Kob, Jian Liu, C. Patrick Royall, Yujie Wang, Hajime Yoshino and Haijun Zhou for discussions. We acknowledge financial support from NSFC (Grants 12161141007, 11935002, 12047503, and 12404290), from Chinese Academy of Sciences (Grant ZDBS-LY-7017 and KGFZD-145-22-13) and from Wenzhou Institute (Grant WIU- CASQD2023009). In this work access was granted to the High-Performance Computing Cluster of Institute of Theoretical Physics - the Chinese Academy of Sciences.
Appendix
Appendix A: Random energy models
1. Model
A REM comprises $2^{N}$ configurations. Each configuration consists of NN Ising spins, whose energy is drawn randomly from a Gaussian probability distribution,
@@ 1. Model
Figure 8: Basin hopping dynamics in MC simulations. Data are obtained for the G-REM at $T=0.75$ with $N=20$. (a) An example of configuration trajectory, where $E(t_{\rm w})$ is the energy of the configuration at $t_{\rm w}$. (b) Corresponding basin trajectory, where $\tilde{E}(t_{\rm w})$ is the energy of the basin or the saddle point at $t_{\rm w}$. With this, a sequence of hopping events with the barrier energy $\Delta E_{i_{k}}$ and the hopping time $\tau_{k}$ are identified. (c) The hopping time probability distribution function (pdf) $\psi(\tau)$ and the conditional distribution $\psi(\tau_{2}|\tau_{1})$ measured in MC simulations, (c) without and (d) with return hops; $\psi^{\rm BT}(\tau)$ is converted from the static distribution $p^{\rm BT}_{\rm G-REM}(\Delta E)$ using the Arrhenius law.
2. Searching for landscape basins and barriers: barrier-tree (BT) algorithm
The BT algorithm searches for all local minima and saddle points in the energy landscape, and organizes them into a barrier tree (see Fig. 2a for an example of a sub-tree). A spin configuration is referred to as a local minimum if its energy is lower than the energy of any adjacent configuration (each configuration has NN adjacent configurations related by a single spin flip). The complete set of local minima are found by exhaustive search. To find the saddle point between two local minima, the algorithm first searches for all possible paths (a path is a series of subsequent spin flips) between the two minima, with the maximum energy point identified along each path – the saddle point is then defined as the lowest energy point among all maxima (min-max).
@@ 2. Searching for landscape basins and barriers: barrier-tree (BT) algorithm
Figure 9: Arrhenius law. (a) Correspondence between $\tau$ and $\Delta E$ in MC simulations of the G-REM at $T=0.75$ with $N=20$, where the color represents the logarithm of the joint probability $\ln\Psi(\tau,\Delta E)$ (see the color bar). (b) The conditional probability $\Psi(\tau|\Delta E=10)$ multiplied by $\tau$; the solid line follows Eq. (10). (b-inset) $A_{0}$ vs NN.
@@
Each glass basin corresponds to a local minimum and a saddle point defined in the above way. A glass basin also contains other configurations that are neither local minima nor saddle points – they form a set of configurations connected to the local minimum whose energies are all below $E_{\rm sp}$. Numerically, we start from the given local minimum, and search for its direct neighbours with a single-spin flip, the neighbours of neighbours, …, until the configuration’s energy $E$ is larger than $E_{\rm sp}$. In this way, for a given realization of the REM with a finite NN, we find all $i=1,2,\ldots,N_{\rm b}$ basins. Each basin corresponds to a set $\mathcal{B}_{i}$ of configurations belonging to it, a local minimum energy $E_{\rm lm}^{i}$, a saddle point energy $E_{\rm sp}^{i}$, and a barrier energy $\Delta E_{i}=E_{\rm sp}^{i}-E_{\rm lm}^{i}$. Note that many configurations do not belong to any basins.
3. Simulations of single-spin flip Metropolis dynamics: Monte Carlo (MC) algorithm
@@
Dynamical trajectories, which are time sequences of configurations $\mathcal{C}(t)$, are obtained by standard single-spin flip MC simulations, starting from random initial configurations. An example is provided in Fig. 8a, where the energy $E(t_{\rm w})$ of the configuration at $t_{\rm w}$ is plotted. In order to obtain the aging function $\Pi_{w}(t_{\rm w})$, our task is to transform the configuration trajectory to a basin trajectory (Fig. 8b). The detailed procedure is described below.
@@
During the dynamics, the system is in the basin $\mathcal{B}_{i}$ if $\mathcal{C}(t)\in\mathcal{B}_{i}$ and it leaves the basin if $\mathcal{C}(t)\notin\mathcal{B}_{i}$ anymore. The duration in the basin defines the hopping time (trapping time) $\tau$. In this way, we identify a sequence of hopping events with $\{\tau_{k},E_{\rm lm}^{i_{k}},E_{\rm sp}^{i_{k}},\Delta E_{i_{k}}\}$, where $k=1,2,\cdots$ and $1\leq i_{k}\leq N_{\rm b}$ denotes that the $k$-th event is in the basin $i_{k}$ (Fig. 8b). Note that the transient time between basins is negligible, i.e., we assume that once the system leaves a basin, it immediately falls into the next basin.
@@
A major computational challenge is the exponential growth of the size $\mathcal{O}(2^{N})$ of the configuration space with the system size NN. For systems with $N>20$ spins, storing the entire landscape would exceed the memory limit of standard computers. To address this issue, we only save the portion of the landscape that is actually explored during the MC dynamical simulations. This trick allows us to simulate up to $N=128$ spins for maximally $10^{9}$ MC steps, with 1 GB memory.
4. Test of the renewal mechanism
Occasionally, after leaving a basin, the system jumps back to the same basin. Following the previous study [38], the two events are merged in such cases. The effects of return jumps are revealed in Fig. 8(c,d) by the conditional probability $p(\tau_{2}|\tau_{1})$, which is the distribution of the next-event hopping time $\tau_{2}$ when the current-event hopping time is $\tau_{1}$. Only when the return jumps are merged, the $p(\tau_{2}|\tau_{1})$ is independent of $\tau_{1}$, which means that the two consecutive jumps are independent. The independence of consecutive jumps (the renewal mechanism) is a basic assumption in the trap model. In order to be consistent with the renewal mechanism, we employ the no-return treatment.
5. Verification of the Arrhenius law
$\tau(\Delta E)=A_{0}\exp(\Delta E/T).$ (9)
We find that the Arrhenius law is consistent with the $\bar{\tau}(\Delta E)$ data ($\bar{\tau}$ is the mean hopping time for the given $\Delta E$) obtained by MC simulations (Fig. 9a), with $A_{0}\sim O(1)$ as shown in Fig. 9b-inset (the time unit is NN).
When NN is finite, for a given $\Delta E$, $\tau$ follows a distribution
$\Psi(\tau|\Delta E)=(1-\eta)^{\tau-1}\eta\approx e^{-\tau\eta}\eta,$ (10)
where $\eta=\exp(-\Delta E/T)$. To derive $\Psi(\tau|\Delta E)$, consider a discretized time $\tau$: if a system is trapped in a basin with a $\Delta E$-barrier for $\tau$ steps, it should remain in the basin for $\tau-1$ steps with a probability $1-\eta$ at each step, and jumps out the basin at the final step with a probability $\eta$. Equation (10) is also verified by the simulation data (see Fig. 9b). Note that the mean hopping time of Eq. (10), $\bar{\tau}(\Delta E)=\int_{0}^{\infty}\tau\Psi(\tau|\Delta E)d\tau$, consistently recovers the Arrhenius law Eq. (9). In short, the Arrhenius law is verified by our MC data.
Appendix B: weak ergodicity breaking order parameter
Figure 10: WEB order parameter $H(w,x)$ by the arcsin law Eq. (2). (a) $H$ versus $w$ for a few different $x$. (b) $H$ versus $x$ for $\omega=0.01,0.1,0.5,2,10,100$ (top to bottom).
Figure 10 shows how $H(w,x)$ depends on $w$ for several given $x$, and on $x$ for several given $w$, as generalization of the theoretical curves in Fig. 1d (blue line) and Fig. 7a.
Appendix C: Weeks-Chandler-Andersen (WCA) model
1. Model
@@
The WCA model is an $80:20$ binary mixture of type A and type B particles interacting with the Lennard-Jones (LJ) potential,
@@
2. Molecular dynamics (MD) simulation method
Simulations of the WCA model are performed at a fixed number density $\rho=1.2$. The MCT temperature at this density is $T_{\rm MCT}\approx 0.28$ [59]. We generate equilibrium configurations at an initial temperature $T_{\rm eq}=5.0$, and then instantly quench the samples to a target temperature $T$. To suppress the wild temperature fluctuations at the initial stage, the aging simulations are performed under the isokinetic ensemble, a variant of the canonical ensemble [79]. The temperature (kinetic energy) is rescaled to the target temperature every 10000 MD steps (100 LJ time units).
We have checked that the aging functions $F_{w}(t_{\rm w})$ obtained by different simulation methods are consistent with each other, apart from the short-time dynamics that are protocol-dependent (see Fig. 11a). Four different simulation methods are compared. (i) MD simulations in the isokinetic ensemble. (ii) MD simulations in the microcanonical (NVE) ensemble with the kinetic energy rescaled to the target value every 50 MD steps (0.1 LJ time units). (iii) MD simulations in the canonical (NVT) ensemble using the Nose-Hoover chain thermostat with a damping parameter of 0.1 LJ time units [79]. (iv) Standard MC simulations in the canonical ensemble. In the MC simulation, we attempt to translate a particle by a displacement vector randomly drawn in a sphere of a radius $5\times 10^{-2}\sigma_{\rm AA}$. The move is accepted according to the Metropolis acceptance rule. In Fig. 11, the MC steps are multiplied by $6.0\times 10^{-4}$ to match the long-time dynamics with those obtained by MD simulations. Further comparison is carried out for the non-equilibrium two-time mean-squared displacement (MSD), $\delta r^{2}(t_{\rm w},t_{\rm w}+t)\equiv\frac{1}{N_{\rm A}}\langle\sum_{i=1}^{N_{\rm A}}|{\bf r}_{i}(t_{\rm w}+t)-{\bf r}_{i}(t_{\rm w})|^{2}\rangle$ (see Fig. 11b). These tests validate the isokinetic ensemble adapted in our MD simulations, which is used to generate the results reported in the current study.
Figure 11: Comparison of the results obtained by different MD and MC algorithms for the WCA model. (a) Aging function $F_{w}(t_{\rm w})$ and (b) non-equilibrium two-time MSD $\delta r^{2}(t_{\rm w},t_{\rm w}+t)$ for $N=1000$ particles at $T=0.25$. In panel (b), $t_{\rm w}\approx 1.0~,16.0,~135,~1.15\times 10^{3},~9.65\times 10^{3},~4.21\times 10^{4}$ from left to right.
Appendix D: Amorphous silica model
1. Model
The atomic interactions in the amorphous silica ($\rm{SiO}_{2}$) is modeled by the Beest Kramer van Santen (BKS) potential, derived from ab initio calculations and lattice dynamics calculations [80]. The BKS potential for both $\rm{Si-O}$ and $\rm{O-O}$ interactions diverges to negative infinity at a small distance $r$, leading to unrealistic attractions at high temperatures. To address this issue, the BKS potential at small distances is replaced by a harmonic repulsive potential, which gives the interactions in the following form:
@@
2. Molecular dynamics simulation method
The MD simulations are carried out using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [81], at a constant volume with a fixed mass density $\rho_{\rm m}=2.36\rm{g/{cm}^{3}}$. The Verlet algorithm with a time step of 1.6 $\rm{fs}$ is used to integrate the equations of motion. We simulate the aging dynamics by quenching the system from a high initial temperature $T_{\rm eq}=8000~{\rm K}$ to a low target temperature $T$. The initial equilibrium configurations at $T_{\rm eq}$ are melted from a crystalline $\rm{SiO}_{2}$ in the NVT ensemble. After quenching, the aging dynamics are simulated with velocities rescaled every 50 steps to keep the target temperature $T$.
3. Estimate of the onset temperature $T_{\rm onset}$
The onset temperature $T_{\rm onset}$ is defined as the characteristic temperature below which glassy dynamics appear. For amorphous silica, $T_{\rm onset}$ has not yet been reported in the literature. In this study two independent methods are employed to estimate $T_{\rm onset}$. The first method is based on the $T$-dependent behavior of the $\alpha$-relaxation time $\tau_{\rm eq}(T)$ in supercooled liquids. One defines $T_{\rm onset}$ at the point where $\tau_{\rm eq}(T)$ departs from the high-$T$ Arrhenius behavior, $\tau_{\rm eq}(T)\sim\exp(E_{\infty}/T)$. Applying this criterion to our simulation data gives $T_{\rm onset}\approx 5800$ K (see Fig. 12a). Note that for amorphous silica, $\tau_{\rm eq}(T)$ is also Arrhenius at low $T$ (with a constant, but larger activation energy than $E_{\infty}$), consistent with the well-know fact that amorphous silica forms strong glasses.
The second method is based on the $T$-dependent behavior of the potential energy $E_{\rm IS}(T)$ of inherent structures [82]. The inherent structures at zero temperature are obtained by minimizing the energy of configurations initially equilibrated at $T$ using the FIRE algorithm [83]. The high-$T$ and low-$T$ data of $E_{\rm IS}(T)$ are fitted by two linear functions, whose intersection defines $T_{\rm onset}\approx 5800$ K (see Fig. 12b). The above two methods give a consistent $T_{\rm onset}$ within the numerical uncertainty.
Figure 12: Onset temperature $T_{\rm onset}$ in the amorphous silica model. (a) $\alpha$-relaxation time $\tau_{\rm eq}$ as a function of $1/T$. The high-$T$ data are fitted to the Arrhenius law (line), and the departing point indicates $T_{\rm onset}$. (b) Inherent structure energy $E_{\rm IS}$ as a function of $T$. The two lines represent linear fitting to low-$T$ and high-$T$ data, whose intersection defines $T_{\rm onset}$. Data are obtained for $N=1536$.
Appendix E: nonergodicity order parameter in structural glasses
Figure 13: Nonergodicity parameter $f$ with respect to the rescaled temperature $T/T_{\rm WEB}$, for WCA and amorphous silica models.
The nonergodicity parameter $f$ is defined by $f=\lim_{t\to\infty}F_{\rm eq}(t)$, where $F_{\rm eq}(t)$ is the equilibrium incoherent scattering function. For the aging data, $f$ corresponds to the plateau value of $F(t_{\rm w},t_{\rm w}+t)$ by sending $t_{\rm w}\to\infty$ first (equilibrium limit) and then $t\to\infty$ (large-time limit); see Fig. 1c. Figure 13 shows the data of $f$ obtained for both WCA and amorphous silica models. Interestingly, the data nearly collapse if plotted as functions of the rescaled temperature $T/T_{\rm WEB}$.
Appendix F: Exponent $\beta$ in structural glasses
According to our theory, the aging function follows logarithmic behavior in the intermediate-time regime, $F_{w}(t_{\rm w})\sim-k\ln t_{\rm w}$. For mean-field models (e.g., REMs), $k\sim 1/N$ (see Eq. 5). For finite-dimensional models, such as the WCA and amorphous silica models, we find that $k\sim N^{-\beta}$ as shown in Fig. 5(c, f) (in the WCA model, $k$ becomes a constant when $N>n$). The $T$-dependent exponent $\beta$ is reported in Fig. 14. For both models, $\beta\approx 0.1-0.4$, which is considerably smaller than the mean-field value $\beta=1$. It is unclear if this discrepancy comes from finite-dimensional effects or pre-asymptotic effects due to insufficiently large systems.
Figure 14: Exponent $\beta$ as a function of $\hat{x}=T/T_{\rm WEB}$ in WCA and amorphous silica models.
@@ 3. Estimate of the onset temperature $T_{\rm onset}$
S1.1 Outline of the general strategy
@@
Here we give a theoretical derivation of the second line ($\tau_{\rm m}<t_{\rm w}<\tau_{\rm th}$) in Eq. (5) of the main text. In Eq. (S5), the normalization constant $\mathcal{C}=1$ in the large-time limit $t_{\rm w}\to\infty$. With $\mathcal{C}=1$, the plateau $H(w,x)$ of the aging function can be determined as given by the arcsin law Eq. (2). In order to obtain the behavior of how the aging function converges to the plateau, it is crucial to consider the next-order $1/t_{\rm w}$-correction to $\mathcal{C}$, as shown below. Note that in this section, we use the BTM distribution Eq. (S3) without finite-size corrections. With Eq. (S3), the average in the denominator of Eq. (S5) becomes,
@@
Figure S1: Comparison between the analytical approximation of the aging function for the BTM and the exact numerical results using the inverse Laplace transform. (a) Large-$t_{\rm w}$ power-law convergence to the asymptotic plateau. (b) Small-$t_{\rm w}$ linear decay. The points are exact results obtained by the numerical inverse Laplace transform of Eqs. (S5) and  (S6). The solid lines are Eq. (S13) in (a) and Eq. (S18) in (b).
@@
where $f(t_{\rm w},\ln\tau)$ is maximized at $\ln\tau^{*}$ with other parameters ($t_{\rm w}$, $T$, etc.) fixed. Equation (S23) is consistent with the numerically evaluated maximum point $\ln\tau^{*}$ (see Fig. S2). As shown by the data, in general $\ln\tau^{*}\sim\ln t_{\rm w}$ in both models (Fig. S2a,b). The GTM data at $x$ and the BTM data at the corresponding $\tilde{x}$ (see Eq. S22) collapse for different $x$ and NN (Fig. S2c,d), and thus Eq. (S23) is verified.
Figure S2: Verification of Eq. (S23). The data points are obtained by numerical evaluation of the maximum $f(t_{\rm w},\ln\tau^{*})$ of the function $f(t_{\rm w},\ln\tau)$ defined in Eq. (S19). As shown by the data, $\ln\tau^{*}$ is linearly proportional to $\ln t_{\rm w}$ in (a) BTM and (b) GTM, for $T=0.21,0.42,0.64,0.85$ (or $x=0.25,0.5,0.75,1.0$), and $N=64$. At the same $T$, the BTM and GTM data do not collapse in (a,b). However, the GTM data (dotted point-line) at $T=T_{\rm c}x$ and the BTM data (solid point-line) at $T=T_{\rm c}\tilde{x}$ collapse, where $\tilde{x}$ is given by Eq. (S22), for different (c) $T$ (with $N=128$ fixed) and (d) NN (with $T=0.42$ fixed).
@@
With Eq. (S21), $\Pi_{w}^{\rm GTM}(t_{\rm w})$ can be conveniently analyzed using the arcsin law, with $x$ in Eq. (S13) replaced by $\tilde{x}$. For large NN, the difference between $\tilde{x}$ and $\hat{x}$, $\tilde{x}-\hat{x}\sim\frac{1}{N}\ln t_{\rm w}$, can be considered as a perturbation to the original form. Expanding around $\hat{x}$ to the first order, we obtain,
@@
where $B=-\frac{T^{2}}{a}\frac{\partial H(w,\hat{x})}{\partial\hat{x}}$. The analytic expression of the logarithmic decay $\frac{B}{N}\ln\left(\frac{t_{\rm w}}{\tau_{0}}\right)$ with $B=-\frac{T^{2}}{a}\frac{\partial H(w,\hat{x})}{\partial\hat{x}}$ is compared to the numerical integration of Eq. (S19) in Fig. S3: the agreement is converged with increasing NN. We further compare the theoretical results $\frac{B}{N}\ln\left(\frac{t_{\rm w}}{\tau_{0}}\right)$ with the MC data in Fig. S4. While the logarithmic form is robust, the coefficient $B$ is not exact due to strong higher-order corrections for small NN in MC simulations. In the main text Fig. 3d, we treat $B$ as a fitting parameter.
@@
Figure S3: Logarithmic decay in the GTM at (a) $T=0.64$ and (b) $T=1$. The data points are obtained by numerical integration of Eq. (S19). The lines are the $\frac{B}{N}\ln\left(\frac{t_{\rm w}}{\tau_{0}}\right)$ term in Eq. (S24).

### v5

@@ Discussion
It is particularly instructive to compare the current GTM with facilitated trap models (FTMs) [79, 80, 81]. The FTM proposed in Refs. [80, 81] describes equilibrium dynamics governed by a distribution of energy barriers (trap energies), $P_{\rm eq}(T,\Delta E)\propto P(\Delta E)\exp(\Delta E/T)$, where $P(\Delta E)$ is the $T$-independent distribution determined by the energy landscape. When a facilitation rule is incorporated into the hopping processes under the constraint of a fixed $P_{\rm eq}(T,\Delta E)$, the trap energies of deep traps effectively diffuse. This diffusion introduces a cutoff time scale $\tau_{p}$ to the trapping time distribution $\psi(\tau)$, even if the original distribution (without facilitation) is unbounded. In contrast, the cutoff time scale $\tau_{\alpha}^{0}(n)$ in the GTM has a fundamentally different, purely thermodynamic origin (the superscript “0” indicates that $\tau_{\alpha}^{0}(n)$ should be estimated without facilitation). It is determined by the typical size $n$ of the activation clusters, corresponding to the mosaic size in the RFOT scenario. Consequently, the actual equilibrium ($\alpha$-) relaxation time $\tau_{\rm eq}$ in systems like the WCA model should be the minimum of $\tau_{p}$ and $\tau_{\alpha}^{0}(n)$; that is, relaxation proceeds via the faster of the two mechanisms (facilitation or RFOT). The influence of facilitation on non-equilibrium aging dynamics, however, remains less clear. In this case, the condition of a fixed $P_{\rm eq}(T,\Delta E)$ no longer applies, and it is unknown whether facilitation would similarly introduce a cutoff time scale through the diffusion of deep trap energies.
@@
Materials and Methods
A. Random energy models
1. Model. A REM comprises $2^{N}$ configurations. Each configuration consists of $N$ Ising spins, whose energy is drawn randomly from a Gaussian probability distribution,
@@ A. Random energy models
2. Searching for landscape basins and barriers: barrier-tree algorithm. The BT algorithm searches for all local minima and saddle points in the energy landscape, and organizes them into a barrier tree (see Fig. 3a for an example of a sub-tree). A spin configuration is referred to as a local minimum if its energy is lower than the energy of any adjacent configuration (each configuration has $N$ adjacent configurations related by a single spin flip). The complete set of local minima are found by exhaustive search. To find the saddle point between two local minima, the algorithm first searches for all possible paths (a path is a series of subsequent spin flips) between the two minima, with the maximum energy point identified along each path – the saddle point is then defined as the lowest energy point among all maxima (min-max).
@@
Each glass basin corresponds to a local minimum and a saddle point defined in the above way. A glass basin also contains other configurations that are neither local minima nor saddle points – they form a set of configurations connected to the local minimum whose energies are all below $E_{\rm sp}$. Numerically, we start from the given local minimum, and search for its direct neighbours with a single-spin flip, the neighbours of neighbours, …, until the configuration’s energy $E$ is larger than $E_{\rm sp}$. In this way, for a given realization of the REM with a finite $N$, we find all $i=1,2,\ldots,N_{\rm b}$ basins. Each basin corresponds to a set $\mathcal{B}_{i}$ of configurations belonging to it, a local minimum energy $E_{\rm lm}^{i}$, a saddle point energy $E_{\rm sp}^{i}$, and a barrier energy $\Delta E_{i}=E_{\rm sp}^{i}-E_{\rm lm}^{i}$. Note that many configurations do not belong to any basins.
@@
3. Simulations of single-spin flip Metropolis dynamics: Monte Carlo algorithm. Dynamical trajectories, which are time sequences of configurations $\mathcal{C}(t)$, are obtained by standard single-spin flip MC simulations, starting from random initial configurations. An example is provided in Fig. S6a, where the energy $E(t_{\rm w})$ of the configuration at $t_{\rm w}$ is plotted. In order to obtain the aging function $\Pi_{w}(t_{\rm w})$, our task is to transform the configuration trajectory to a basin trajectory (Fig. S6b). The detailed procedure is described below.
@@
During the dynamics, the system is in the basin $\mathcal{B}_{i}$ if $\mathcal{C}(t)\in\mathcal{B}_{i}$ and it leaves the basin if $\mathcal{C}(t)\notin\mathcal{B}_{i}$ anymore. The duration in the basin defines the hopping time (trapping time) $\tau$. In this way, we identify a sequence of hopping events with $\{\tau_{k},E_{\rm lm}^{i_{k}},E_{\rm sp}^{i_{k}},\Delta E_{i_{k}}\}$, where $k=1,2,\cdots$ and $1\leq i_{k}\leq N_{\rm b}$ denotes that the $k$-th event is in the basin $i_{k}$. Note that the transient time between basins is negligible, i.e., we assume that once the system leaves a basin, it immediately falls into the next basin.
@@
A major computational challenge is the exponential growth of the size $\mathcal{O}(2^{N})$ of the configuration space with the system size $N$. For systems with $N>20$ spins, storing the entire landscape would exceed the memory limit of standard computers. To address this issue, we only save the portion of the landscape that is actually explored during the MC dynamical simulations. This trick allows us to simulate up to $N=128$ spins for maximally $10^{9}$ MC steps, with 1 GB memory.
B. Weeks-Chandler-Andersen model
@@
1. Model. The WCA model is an $80:20$ binary mixture of type A and type B particles interacting with the Lennard-Jones (LJ) potential,
@@
2. Molecular dynamics simulation method. Simulations of the WCA model are performed at a fixed number density $\rho=1.2$. The MCT temperature at this density is $T_{\rm MCT}\approx 0.28$ [59]. We generate equilibrium configurations at an initial temperature $T_{\rm eq}=5.0$, and then instantly quench the samples to a target temperature $T$. To suppress the wild temperature fluctuations at the initial stage, the aging simulations are performed under the isokinetic ensemble, a variant of the canonical ensemble [85]. The temperature (kinetic energy) is rescaled to the target temperature every 10000 MD steps (100 LJ time units).
C. Amorphous silica model
1. Model. The atomic interactions in the amorphous silica ($\rm{SiO}_{2}$) is modeled by the Beest Kramer van Santen (BKS) potential, derived from ab initio calculations and lattice dynamics calculations [86]. The BKS potential for both $\rm{Si-O}$ and $\rm{O-O}$ interactions diverges to negative infinity at a small distance $r$, leading to unrealistic attractions at high temperatures. To address this issue, the BKS potential at small distances is replaced by a harmonic repulsive potential, which gives the interactions in the following form:
@@ C. Amorphous silica model
2. Molecular dynamics simulation method. The MD simulations are carried out using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [87], at a constant volume with a fixed mass density $\rho_{\rm m}=2.36\rm{g/{cm}^{3}}$. The Verlet algorithm with a time step of 1.6 $\rm{fs}$ is used to integrate the equations of motion. We simulate the aging dynamics by quenching the system from a high initial temperature $T_{\rm eq}=8000~{\rm K}$ to a low target temperature $T$. The initial equilibrium configurations at $T_{\rm eq}$ are melted from a crystalline $\rm{SiO}_{2}$ in the NVT ensemble. After quenching, the aging dynamics are simulated with velocities rescaled every 50 steps to keep the target temperature $T$.
Acknowledgments
We warmly thank Wenyue Fan, Gang Huang, Wencheng Ji, Walter Kob, Jian Liu, C. Patrick Royall, Yujie Wang, Hajime Yoshino and Haijun Zhou for discussions.
Funding: Y.J. acknowledges financial support from NSFC (Grants 12161141007, 11935002 and 12047503), from Chinese Academy of Sciences (Grant ZDBS-LY-7017 and KGFZD-145-22-13) and from Wenzhou Institute (Grant WIU- CASQD2023009). D.P. acknowledges support from NSFC (Grants 12404290). In this work access was granted to the High-Performance Computing Cluster of Institute of Theoretical Physics - the Chinese Academy of Sciences.
Author Contributions: All authors contributed equally to this work.
Competing Interests: The authors declare that they have no competing interests.
Data and Materials Availability: All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials. The simulation program is available at https://doi.org/10.5281/zenodo.17737463.
@@ Acknowledgments
S1.1 Outline of the general strategy.
@@
FIG. S1: WEB order parameter $H(w,x)$ by the arcsin law Eq. (2). (a) $H$ versus $w$ for a few different $x$. (b) $H$ versus $x$ for $\omega=0.01,0.1,0.5,2,10,100$ (top to bottom).
@@
Here we give a theoretical derivation of the second line ($\tau_{\rm m}<t_{\rm w}<\tau_{\rm th}$) in Eq. (5) of the main text. In Eq. (S5), the normalization constant $\mathcal{C}=1$ in the large-time limit $t_{\rm w}\to\infty$. With $\mathcal{C}=1$, the plateau $H(w,x)$ of the aging function can be determined as given by the arcsin law Eq. (2) (see Fig. S1). In order to obtain the behavior of how the aging function converges to the plateau, it is crucial to consider the next-order $1/t_{\rm w}$-correction to $\mathcal{C}$, as shown below. Note that in this section, we use the BTM distribution Eq. (S3) without finite-size corrections. With Eq. (S3), the average in the denominator of Eq. (S5) becomes,
@@
FIG. S2: Comparison between the analytical approximation of the aging function for the BTM and the exact numerical results using the inverse Laplace transform. (a) Large-$t_{\rm w}$ power-law convergence to the asymptotic plateau. (b) Small-$t_{\rm w}$ linear decay. The points are exact results obtained by the numerical inverse Laplace transform of Eqs. (S5) and  (S6). The solid lines are Eq. (S13) in (a) and Eq. (S18) in (b).
@@
where $f(t_{\rm w},\ln\tau)$ is maximized at $\ln\tau^{*}$ with other parameters ($t_{\rm w}$, $T$, etc.) fixed. Equation (S23) is consistent with the numerically evaluated maximum point $\ln\tau^{*}$ (see Fig. S3). As shown by the data, in general $\ln\tau^{*}\sim\ln t_{\rm w}$ in both models (Fig. S3a,b). The GTM data at $x$ and the BTM data at the corresponding $\tilde{x}$ (see Eq. S22) collapse for different $x$ and $N$ (Fig. S3c,d), and thus Eq. (S23) is verified.
FIG. S3: Verification of Eq. (S23). The data points are obtained by numerical evaluation of the maximum $f(t_{\rm w},\ln\tau^{*})$ of the function $f(t_{\rm w},\ln\tau)$ defined in Eq. (S19). As shown by the data, $\ln\tau^{*}$ is linearly proportional to $\ln t_{\rm w}$ in (a) BTM and (b) GTM, for $T=0.21,0.42,0.64,0.85$ (or $x=0.25,0.5,0.75,1.0$), and $N=64$. At the same $T$, the BTM and GTM data do not collapse in (a,b). However, the GTM data (dotted point-line) at $T=T_{\rm c}x$ and the BTM data (solid point-line) at $T=T_{\rm c}\tilde{x}$ collapse, where $\tilde{x}$ is given by Eq. (S22), for different (c) $T$ (with $N=128$ fixed) and (d) NN (with $T=0.42$ fixed).
@@
With Eq. (S21), $\Pi_{w}^{\rm GTM}(t_{\rm w})$ can be conveniently analyzed using the arcsin law, with $x$ in Eq. (S13) replaced by $\tilde{x}$. For large $N$, the difference between $\tilde{x}$ and $\hat{x}$, $\tilde{x}-\hat{x}\sim\frac{1}{N}\ln t_{\rm w}$, can be considered as a perturbation to the original form. Expanding around $\hat{x}$ to the first order, we obtain,
@@
where $B=-\frac{T^{2}}{a}\frac{\partial H(w,\hat{x})}{\partial\hat{x}}$. The analytic expression of the logarithmic decay $\frac{B}{N}\ln\left(\frac{t_{\rm w}}{\tau_{0}}\right)$ with $B=-\frac{T^{2}}{a}\frac{\partial H(w,\hat{x})}{\partial\hat{x}}$ is compared to the numerical integration of Eq. (S19) in Fig. S4: the agreement is converged with increasing $N$. We further compare the theoretical results $\frac{B}{N}\ln\left(\frac{t_{\rm w}}{\tau_{0}}\right)$ with the MC data in Fig. S5. While the logarithmic form is robust, the coefficient $B$ is not exact due to strong higher-order corrections for small $N$ in MC simulations. In the main text Fig. 3d, we treat $B$ as a fitting parameter.
@@
FIG. S4: Logarithmic decay in the GTM at (a) $T=0.64$ and (b) $T=1$. The data points are obtained by numerical integration of Eq. (S19). The lines are the $\frac{B}{N}\ln\left(\frac{t_{\rm w}}{\tau_{0}}\right)$ term in Eq. (S24).